Newly purchased tires of a particular type are supposed to be filled to a pressure of 30 psi. Let μ denote the true average pressure. A test is to be carried out to decide whether μ differs from the target value. Determine the P-value for each of the following z test statistic values. (Round your answers to four decimal places.)

(a)

2.19

(b)

−1.72

(c)

−0.52

(d)

1.47

(e)

−5.05
Answer #1

Solution:

This is a two tailed test,

a)

P- Value = 2*P(Z > z )

= 2 * ( 1 - P(Z < 2.19 ))

= 2 * 0.0143

= 0.0286

b)

P- Value = 2*P(Z < z )

= 2 * P(Z < -1.72 )

= 2 * 0.0427

= 0.0854

c)

P- Value = 2*P(Z < z )

= 2 * P(Z < -0.52 )

= 2 * 0.3015

= 0.6030

d)

P- Value = 2*P(Z > z )

= 2 * ( 1 - P(Z < 1.47 ))

= 2 * 0.0708

= 0.1416

e)

P- Value = 2*P(Z < z )

= 2 * P(Z < -5.05 )

= 2 * 0.000

= 0.0000